A consumer has $25 in budget to purchase goods X & Y. Assume Px = 3 and Py=1. What is the maximum amount of Y that she can buy?


1
Expert's answer
2021-02-18T07:13:03-0500

Solution:

Derive the income function (budget line):

I = PxX + PyY

25 = 3X + Y

Y = 25 – 3X

The optimal consumption bundle is where the slope of the indifference curve"(\frac{MUx}{MUy} )" is equal to the slope of the budget line "(\frac{Px}{Py} )" in absolute value terms.

MUx = Y and MUy = X, therefore "(\frac{MUx}{MUy} )" = "(\frac{Y}{X} )"


"(\frac{Px}{Py} )" = "(\frac{3}{1} )" = 3, Therefore, "(\frac{Y}{X} )" = 3 or X = "(\frac{Y}{3} )"


Substitute this into the budget line to get:

Y = 25 – 3X

Y = 25 – 3 "(\frac{Y}{3} )"

Y = 25 – Y

Y + Y = 25

2Y = 25

Y = "(\frac{25}{2} )" = 13

To get X:

X = "(\frac{Y}{3} )" = "(\frac{13}{3} )" = 4


Therefore, the optimum consumption bundle (x,y) = (4, 13)


The maximum amount of Y she can buy is 13 units